# Zhang, Liu, and Zhao

### v3

@@ 10.4 Verification of Assumption 1
We next show that for any $\mathcal{C}\in\mathfrak{C}$ with $\mathcal{S}^{d}\bigcap\mathcal{C}=\emptyset$, given $\varepsilon>0$ and $\rho>0$, there exists $\widehat{\lambda}_{q}>0$ such that $r^{\varepsilon,\rho}(\theta)>\widehat{\lambda}_{q}$ for any $\theta\in\mathcal{C}$. For such $\mathcal{C}\in\mathfrak{C}$, we have $\mathcal{S}^{*}_{\mathcal{C}}\subseteq\partial{(\mathcal{C})}$ and $\mathcal{S}^{*}_{\mathcal{C}}\bigcap\mathcal{S}^{d}=\emptyset$. Since $\mathcal{S}^{d}\bigcap\mathrm{int}(\mathcal{C})=\emptyset$, by arguments above there exists a polytope $\mathcal{C}^{\prime}\in\mathfrak{C}$ such that $\mathcal{S}^{*}_{\mathcal{C}}\subseteq\mathcal{C}^{\prime}$ and $\min_{\theta\in\mathcal{S}^{*}_{\mathcal{C}}}\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}^{\prime}})>0$. Let $I_{\mathcal{C}^{\prime}}$ denote the index mapping of $f$ corresponding to $\mathcal{C}^{\prime}$. Due to the Lipschitz continuity of $F$ on $\Theta$, for any $\varepsilon>0$ there exist $\delta>0$ and $\Delta>0$ such that any $\theta\in\mathcal{C}$ with $\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})\leq\delta$ satisfies $I_{\mathcal{C}^{\prime}}\in\mathcal{I}^{\varepsilon}(\theta)$ and $F(\theta)>\min_{\theta^{\prime}\in\mathcal{C}^{\prime}}F(\theta^{\prime})+\Delta$; consequently, for any given $\rho>0$, there exists $\lambda_{1}>0$ such that $r^{\varepsilon,\rho}(\theta)>\lambda_{1}$. On the other hand, for any point $\theta\in\mathcal{C}$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})>\delta$, by the first statement of Proposition 10.6, there exists $\lambda_{2}>0$ such that $r^{\varepsilon,\rho}(\theta)>\lambda_{2}$. Combining the above two cases, we obtain the desired statement. It yields that there exists $\lambda_{q}>0$ such that for any $\theta\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\leq{\lambda}_{q}$, we have $\mathcal{S}^{d}\bigcap\mathcal{C}\neq\emptyset$ and $\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}=\mathcal{S}^{*}_{\mathcal{C}}$ for any $\mathcal{C}\in\mathfrak{C}_{\theta}\coloneqq\{\mathcal{C}^{\prime}\in\mathfrak{C}\mid\theta\in\mathcal{C}^{\prime}\}$.
Now, consider a point $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$ satisfying $r^{\varepsilon,\rho}(\theta)\leq{\lambda}_{q}$. Given a polytope $\mathcal{C}\in\mathfrak{C}_{\theta}$ denoted by $\mathcal{C}\coloneqq\{\theta\in\mathbb{R}^{q}\mid B\theta\leq b\}$, since $\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}=\mathcal{S}^{*}_{\mathcal{C}}$, we also have $\theta\in\mathcal{C}\backslash\mathcal{S}^{*}_{\mathcal{C}}$. With $f(\theta)\coloneqq A\theta$ over $\mathcal{C}$ for some matrix $A$ and $a^{*}\coloneqq A\theta^{*}$ for $\theta^{*}\in\mathcal{S}^{*}_{\mathcal{C}}$, we obtain $\mathcal{S}^{*}_{\mathcal{C}}=\{\theta\in\mathbb{R}^{q}\mid B\theta\leq b,A\theta\leq a^{*},-A\theta\leq-a^{*}\}$. By compactness of $\mathcal{C}$ and Hoffman’s lemma (Hoffman 1952), there exists $c>0$ depending on $\mathcal{C}$ such that for any $\theta\in\mathcal{C}$, there exists $\theta^{d}\in\mathcal{S}^{*}_{\mathcal{C}}=\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}$ satisfying
@@
11 Analysis of PADR-based ERM for Constrained SP
@@ 11.1 Asymptotic Consistency of Constrained PADR-based ERM
  1. (${}^{\prime}$)
    There exists a continuous function $f^{*}_{\gamma_{0}}$ that is optimal to problem (11.1).
By Assumptions (A2) and (${}^{\prime}$), $f^{*}_{\gamma_{0}}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$ for some $L_{z}^{\prime}>0$ and $M_{z}^{\prime}>0$. We denote by ~${\widetilde{\mathcal{H}}}^{K}_{\mu,\gamma}\coloneqq{\mathcal{H}}_{\mu}^{K}\bigcap\left\{f\in\mathcal{F}\mid\psi_{j}(f(x^{s}))+\gamma\leq 0,\,s\in[n],\,j\in[J]\right\}$ and $f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ the feasible PADR class and solution to the constrained ERM problem (5) respectively. Then asymptotic consistency of the constrained PADR-based ERM model is presented below.
Theorem 11.1 (Asymptotic Consistency of Constrained PADR-based ERM)
Suppose that Assumptions (A1)-(A3) and (A5) hold, and that Assumption (${}^{\prime}$) holds for some $\gamma_{0}>\Gamma\coloneqq 2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\psi}L_{z}^{\prime}\bar{X}K^{-1/p}$. With $\gamma\in(0,\gamma_{0}-\Gamma)$, $\mu\geq\max\{\frac{1}{2}L_{z}^{\prime}K^{1/p},\frac{1}{4}pL_{z}^{\prime}\bar{X}K^{1/p}+\frac{1}{2}M_{z}^{\prime}\}$, and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $S$,
$R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})\leq$ $R(f^{*}_{\gamma_{0}})+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}^{\prime}\bar{X}K^{-1/p}$
@@
Furthermore, with $\mu=\mathcal{O}(K^{1/p})$, $\gamma=\mathcal{O}(K^{-1/p})$, $\gamma_{0}=\mathcal{O}(K^{-1/p})$, and $K=\mathcal{O}(n^{p/(p+4)})$, suppose that there exists $\bar{\gamma}_{0}>0$ such that for any $\gamma\in[0,\bar{\gamma}_{0}]$, Assumption (${}^{\prime}$) holds and $f^{*}_{\gamma}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$. Then $R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})$ converges to the optimal cost of (5) in probability, as $n$ goes to infinity.
Proof 11.2
@@
$R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$ $=R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{S}(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{S}(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
$\leq R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{S}(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{S}(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
$\leq\underbrace{2\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{S}(f)\right\rvert}_{\mathcal{E}_{1}}+\underbrace{R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})}_{\mathcal{E}_{2}},$
where the last inequality holds since ~$\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}\subseteq{\mathcal{H}}_{\mu}^{K}$. The first term $\mathcal{E}_{1}$ can be directly bounded by Proposition 2.2, that is, given $\delta>0$, with probability at least $1-\delta$ over the draw of $S$,
@@
By combining bounds of $\mathcal{E}_{1}$, $\mathcal{E}_{21}$, and $\mathcal{E}_{22}$, we derive the desired inequality in the theorem, and asymptotic consistency follows instantly.∎
11.2 Convergence of Enhanced SMM to B-stationarity
According to (Cui and Pang 2021), for a constrained problem with a B-differentiable objective, a B-stationary point is defined as follows.
@@ 11.2 Convergence of Enhanced SMM to B-stationarity
We say that $\theta^{b}$ is a B-stationary point of the objective $F(\cdot)$ on a closed set $\bar{\Theta}\subseteq\Theta$ if $\theta^{b}\in\bar{\Theta}$ and $F^{\prime}(\theta^{b};v)\geq 0$ for any $v\in\mathcal{T}(\theta^{b};\bar{\Theta})$.
The B-stationarity coincides with d-stationarity if $\bar{\Theta}$ is convex. Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ in problem (5) are B-differentiable. With $\bar{\Theta}=\left\{\theta\in\Theta\mid G_{\gamma}(\theta)\leq 0\right\}$, we denote the set of B-stationary points of (5) by $\mathcal{S}^{b}$.
With Assumptions 1-1 imposed in the context of the penalized problem (5), we can apply the enhanced SMM algorithm to solve (5) and obtain the same convergence results in terms of d-stationarity as in Theorems 4.4 and 4.5. To further obtain the convergence to B-stationarity, we intend to connect d-stationarity of (5) with B-stationarity of (5) by applying the following exact penalty result from Cui and Pang (2021). Note that by the finiteness of $S$ and compactness of $\Theta$, for any $\xi\in S$, $f(\cdot;\xi)$ is Lipschitz continuous with modulus $L_{f}\coloneqq 2\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$ and uniformly bounded with $\left\lVert f(\cdot;\xi)\right\rVert_{\infty}\leq M_{f}\coloneqq L_{f}\max_{\theta\in\Theta}\left\lVert\theta\right\rVert$, and $\varphi(\cdot;\xi)$ is Lipschitz continuous on $[-M_{f},M_{f}]^{d}$ with uniform modulus denoted by $L^{\prime}_{\varphi}$.
Lemma 11.4 (Cui and Pang (2021), Proposition 9.2.2)
@@
then with $\lambda>L^{\prime}_{\varphi}L_{f}/l_{c}$, any d-stationary point of (5) is feasible to (5) and thus is a B-stationary point of (5).
Condition (24) requires that the maximum descent rates of $G_{\gamma}$ at points infeasible to (5) are bounded away from 0. Thus by the Lipschitz continuity of $F$, at any infeasible point, we can always find a descent direction in the penalized problem (5) with a sufficiently large (but still finite) penalty parameter $\lambda$, ensuring that the point can not be a d-stationary point of (5). Nonetheless, it is not easy to verify the technical condition (24) when $G_{\gamma}$ is in a composite structure. The following assumption on $\{\psi_{j}\}_{j\in[J]}$ provides a sufficient condition for (24):
@@
Suppose that Assumption (B1) holds, and that $\Theta\subseteq\mathbb{R}^{q}$ is a compact convex set with $0\in\Theta$ and $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$. Then condition (24) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$. Moreover, with $\lambda>\bar{\lambda}\coloneqq\frac{n\sqrt{q}\mu L^{\prime}_{\varphi}L_{f}}{c_{\psi}l_{\psi}}$, any d-stationary point of (5) is a B-stationary point of problem (5).
Proof 11.6
Proof of Proposition 11.5. By Lemma 11.4, it suffices to show that condition (24) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$. For any $\xi\in S$ and $j\in[J]$, denote $G_{\gamma}(\theta;\xi,j)\coloneqq\max\left\{\psi_{j}(f(\theta;\xi))+\gamma,0\right\}$. For a point $\bar{\theta}\in\Theta$ infeasible to the constraint $G_{\gamma}(\theta)\coloneqq\mathbb{E}\left[\sum_{j=1}^{J}G_{\gamma}(\theta;\widetilde{\xi}_{n},j)\right]\leq 0$, there exist $\bar{j}\in[J]$ and $\bar{\xi}\in S$ such that $\psi_{\bar{j}}(f(\bar{\theta};\bar{\xi}))+\gamma>0$.
For any $\xi\in S$, we denote $\bar{z}=f(\bar{\theta};\xi)$ and $\bar{v}=-\frac{\bar{\theta}}{\left\lVert\bar{\theta}\right\rVert}$ for simplicity. Since $f(\cdot;\xi)$ is piecewise linear, there exists $A\in\mathbb{R}^{d\times q}$ dependent on $\xi$ and $\bar{\theta}$ such that $f(\bar{\theta};\xi)=A\bar{\theta}$ and $f(\cdot;\xi)^{\prime}(\bar{\theta};\bar{v})=A\bar{v}$. Let
@@
For $j\in\mathcal{A}_{<}$, it is clear that $G_{\gamma}(\cdot;\xi,j)^{\prime}(\bar{\theta};\bar{v})=0$. By Assumption (B1), we have that for $j\in\mathcal{A}_{>}$,
$G_{\gamma}(\cdot;\xi,j)^{\prime}(\bar{\theta};\bar{v})=\psi_{j}^{\prime}(f(\bar{\theta};\xi);f(\cdot;\xi)^{\prime}(\bar{\theta};\bar{v}))=\psi_{j}^{\prime}\left(\bar{z};-\frac{\bar{z}}{\left\lVert\bar{\theta}\right\rVert}\right)\leq-\frac{c_{\psi}l_{\psi}}{\left\lVert\bar{\theta}\right\rVert}\leq-\frac{c_{\psi}l_{\psi}}{\sqrt{q}\mu},$
@@
$\min_{v\in\mathcal{T}(\bar{\theta};\Theta);\left\lVert v\right\rVert\leq 1}G_{\gamma}^{\prime}(\bar{\theta};v)\leq G_{\gamma}^{\prime}(\bar{\theta};\bar{v})\leq\frac{1}{n}G_{\gamma}(\cdot;\bar{\xi},\bar{j})^{\prime}(\bar{\theta};\bar{v})\leq-\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu},$
@@
Finally, from Proposition 11.5 and convergence results in Section 4, we summarize the convergence properties of the enhanced SMM with exact penalization in the theorem below, where Assumptions 1 and 2 are adjusted to surrogate families of $F(\cdot;\xi)$ and $\psi_{j}(f(\cdot;\xi))$ for any $\xi\in S$ and $j\in[J]$, and Assumption 1 is adjusted to $V(\cdot;\lambda)$. The proof is similar to those of Theorems 4.4 and 4.5 and thus is omitted here.
Theorem 11.7 (Convergence of Enhanced SMM with exact penalization)
Under Assumption 1 and the same settings of Proposition 11.5, for problem (5) with $\lambda\geq\bar{\lambda}$, let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon\geq 0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$. The following statements hold:

### v4

@@ 8.5 Verification of Assumption (B6)
We next show that for any $\mathcal{C}\in\mathfrak{C}$ with $\mathcal{S}^{d}\bigcap\mathcal{C}=\emptyset$, given $\varepsilon>0$ and $\rho>0$, there exists $\widehat{\lambda}_{q}>0$ such that $r^{\varepsilon,\rho}(\theta)>\widehat{\lambda}_{q}$ for any $\theta\in\mathcal{C}$. For such $\mathcal{C}\in\mathfrak{C}$, we have $\mathcal{S}^{*}_{\mathcal{C}}\subseteq\partial{(\mathcal{C})}$ and $\mathcal{S}^{*}_{\mathcal{C}}\bigcap\mathcal{S}^{d}=\emptyset$. Since $\mathcal{S}^{d}\bigcap\mathrm{int}(\mathcal{C})=\emptyset$, by arguments above there exists a polytope $\mathcal{C}^{\prime}\in\mathfrak{C}$ such that $\mathcal{S}^{*}_{\mathcal{C}}\subseteq\mathcal{C}^{\prime}$ and $\min_{\theta\in\mathcal{S}^{*}_{\mathcal{C}}}\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}^{\prime}})>0$. Let $\mathbf{I}_{\mathcal{C}^{\prime}}$ denote the index combination of $f$ corresponding to $\mathcal{C}^{\prime}$. Due to the Lipschitz continuity of $F$ on $\Theta$, for any $\varepsilon>0$ there exist $\delta>0$ and $\Delta>0$ such that any $\theta\in\mathcal{C}$ with $\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})\leq\delta$ satisfies $\mathbf{I}_{\mathcal{C}^{\prime}}\in\mathcal{I}^{\varepsilon}(\theta)$ and $F(\theta)>\min_{\theta^{\prime}\in\mathcal{C}^{\prime}}F(\theta^{\prime})+\Delta$; consequently, for any given $\rho>0$, there exists $\lambda_{1}>0$ such that $r^{\varepsilon,\rho}(\theta)>\lambda_{1}$. On the other hand, for any point $\theta\in\mathcal{C}$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})>\delta$, by the first statement of Proposition 8.6, there exists $\lambda_{2}>0$ such that $r^{\varepsilon,\rho}(\theta)>\lambda_{2}$. Combining the above two cases, we obtain the desired statement. It yields that there exists $\lambda_{q}>0$ such that for any $\theta\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\leq{\lambda}_{q}$, we have $\mathcal{S}^{d}\bigcap\mathcal{C}\neq\emptyset$ and $\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}=\mathcal{S}^{*}_{\mathcal{C}}$ for any $\mathcal{C}\in\mathfrak{C}_{\theta}\coloneqq\{\mathcal{C}^{\prime}\in\mathfrak{C}\mid\theta\in\mathcal{C}^{\prime}\}$.
[…]
@@
9 Constrained SP with Covariate Information
In this section, we study problem (1) with deterministic constraints $\psi_{j}(z)\leq 0$ for $j\in[J]$. We first provide the formulation of the PADR-based ERM method. Then, we establish the asymptotic consistency of the constrained PADR-based ERM model in Section 9.1 and develop ESMM algorithm with exact penalization in Section 9.2. We present the numerical results on the Newsvendor problem with convex and nonconvex constraints in Section 9.3.
[…]
  1. (C1)
    Finite scenarios: when the support set $\mathcal{X}$ of the feature vector $x$ contains a finite number of known scenarios, feasibility can be guaranteed by imposing constraints on the decision rule under all scenarios.
  2. (C2)
    Box constraints: when deterministic constraints lead to a box-constrained set $\{z\in\mathbb{R}^{d}:\left\lVert z\right\rVert_{\infty}\leq M\}$, a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ can be used, where ${\mathcal{H}}_{\mu}^{K}$ represents the PA hypothesis class, and the max/min are element-wise operators. This ensures feasibility of the ERM solution.
Following the excess risk decomposition in Section 2.2, it is straightforward to see that the PADR-based ERM model for the three special cases (C1) and (C2) is nonasymptotically consistent with the original SP problem (9).
[…]
We next show that the PADR-based ERM model (9) is asymptotically consistent with (9).
@@ 9.1 Asymptotic consistency of constrained PADR-based ERM
  1. (A4)
    Given $\gamma_{0}>0$, there exists a Lipschitz continuous function $f^{*}_{\gamma_{0}}$ that is optimal to problem (9.1).
By Assumptions (A2) and (A4), $f^{*}_{\gamma_{0}}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$ for some $L_{z}^{\prime}>0$ and $M_{z}^{\prime}>0$. We first provide the excess risk bound of the restricted ERM model (9) to the restricted SP problem (9.1).
Proposition 9.1 (Excess risk bound of the restricted constrained ERM)
Suppose that Assumptions (A1)-(A3) and (A5) hold, and that Assumption (A4) holds for some $\gamma_{0}>\Gamma\coloneqq 2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\psi}L_{z}^{\prime}\bar{X}K^{-1/p}$. With $\gamma\in(0,\gamma_{0}-\Gamma)$, $\mu\geq\max\{\frac{1}{2}L_{z}^{\prime}K^{1/p},\frac{1}{4}pL_{z}^{\prime}\bar{X}K^{1/p}+\frac{1}{2}M_{z}^{\prime}\}$, and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
$R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})\leq$ $R(f^{*}_{\gamma_{0}})+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}^{\prime}\bar{X}K^{-1/p}$
@@
Proof 9.2
@@
$R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$ $=R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{\Xi}(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{\Xi}(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
$\leq R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{\Xi}(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{\Xi}(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
$\leq\underbrace{2\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert}_{\mathcal{E}_{1}}+\underbrace{R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})}_{\mathcal{E}_{2}},$
where the last inequality holds since ~$\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}\subseteq{\mathcal{H}}_{\mu}^{K}$. The first term $\mathcal{E}_{1}$ can be directly bounded by Proposition 2.2, that is, given $\delta>0$, with probability at least $1-\delta$ over the draw of $\Xi$,
@@
By combining bounds of $\mathcal{E}_{1}$, $\mathcal{E}_{21}$, and $\mathcal{E}_{22}$, we derive the desired bound.∎
With the excess risk bound in Proposition 9.1, and the probabilistic feasibility guarantee according to Luedtke and Ahmed (2008, Theorem 10), we obtain the asymptotic consistency of restricted PADR-based ERM model to the original constrained SP (9).
Theorem 9.3 (Asymptotic consistency of constrained PADR-based ERM)
Suppose that Assumptions (A1)-(A3) and (A5) hold, and that $f^{*}_{\gamma}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$ for any $\gamma\in[0,\bar{\gamma}]$ with some $\bar{\gamma}>0$. Then for any $n>0$, there exist $\delta_{n}>0$ and $\epsilon_{n}>0$ such that $f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ satisfies $\mathbb{P}_{X}\big{\{}\psi_{j}(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}(X))\leq 0,\,j\in[J]\big{\}}\geq 1-\epsilon_{n}$ with probability $1-\delta_{n}$, where $\epsilon_{n}$ and $\delta_{n}$ converge to 0 as $n$ goes to infinity. Moreover, with $K=\mathcal{O}(n^{p/(p+4)})$, $\mu=\mathcal{O}(K^{1/p})$, and $\gamma=\mathcal{O}(K^{-1/p})$, $R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})$ converges to the optimal cost of (9) in probability, as $n$ goes to infinity.
9.2 Convergence of ESMM to B-stationarity
[…]
@@ 9.2 Convergence of ESMM to B-stationarity
We say that $\theta^{b}$ is a B-stationary point of the objective $F(\cdot)$ on a closed set $\bar{\Theta}\subseteq\Theta$ if $\theta^{b}\in\bar{\Theta}$ and $F^{\prime}(\theta^{b};v)\geq 0$ for any $v\in\mathcal{T}(\theta^{b};\bar{\Theta})$. where $\mathcal{T}(\theta^{b};\bar{\Theta})$ is the tangent cone of $\bar{\Theta}$ at $\theta^{b}$.
The B-stationarity coincides with d-stationarity if $\bar{\Theta}$ is convex. Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ in problem (9.2) are B-differentiable. With $\bar{\Theta}=\left\{\theta\in\Theta\mid G_{\gamma}(\theta)\leq 0\right\}$, we denote the set of B-stationary points of (9.2) by $\mathcal{S}^{b}$.
[…]
@@
then with $\lambda>L^{\prime}_{\varphi}L_{f}/l_{c}$, any d-stationary point of (9.2) is feasible to (9.2) and thus is a B-stationary point of (9.2).
Condition (32) requires that the maximum descent rates of $G_{\gamma}$ at points infeasible to (9.2) are bounded away from 0. Thus by the Lipschitz continuity of $F$, at any infeasible point, we can always find a descent direction in the penalized problem (9.2) with a sufficiently large (but still finite) penalty parameter $\lambda$, ensuring that the point can not be a d-stationary point of (9.2). Nonetheless, it is not easy to verify the technical condition (32) when $G_{\gamma}$ is in a composite structure. The following assumption on $\{\psi_{j}\}_{j\in[J]}$ provides a sufficient condition for (32):
@@
Suppose that Assumption (B7) holds, and that $\Theta\subseteq\mathbb{R}^{q}$ is a compact convex set with $0\in\Theta$ and $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$. Then condition (32) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$. Moreover, with $\lambda>\bar{\lambda}\coloneqq\frac{n\sqrt{q}\mu L^{\prime}_{\varphi}L_{f}}{c_{\psi}l_{\psi}}$, any d-stationary point of (9.2) is a B-stationary point of problem (9.2).
Proof 9.7
Proof of Proposition 9.6. By Lemma 9.5, it suffices to show that condition (32) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$. For any $\xi\in\Xi$ and $j\in[J]$, denote $G_{\gamma}(\theta,\xi;j)\coloneqq\max\left\{\psi_{j}(f(\theta,\xi))+\gamma,0\right\}$. For a point $\bar{\theta}\in\Theta$ infeasible to the constraint $G_{\gamma}(\theta)\coloneqq\mathbb{E}\left[\sum_{j=1}^{J}G_{\gamma}(\theta,\widetilde{\xi}_{n};j)\right]\leq 0$, there exist $\bar{j}\in[J]$ and $\bar{\xi}\in\Xi$ such that $\psi_{\bar{j}}(f(\bar{\theta},\bar{\xi}))+\gamma>0$.
For any $\xi\in\Xi$, we denote $\bar{z}=f(\bar{\theta},\xi)$ and $\bar{v}=-\frac{\bar{\theta}}{\left\lVert\bar{\theta}\right\rVert}$ for simplicity. Since $f(\cdot,\xi)$ is piecewise linear, there exists $A\in\mathbb{R}^{d\times q}$ dependent on $\xi$ and $\bar{\theta}$ such that $f(\bar{\theta},\xi)=A\bar{\theta}$ and $f(\cdot,\xi)^{\prime}(\bar{\theta};\bar{v})=A\bar{v}$. Let
@@
For $j\in\mathcal{A}_{<}$, it is clear that $G_{\gamma}(\cdot,\xi;j)^{\prime}(\bar{\theta};\bar{v})=0$. By Assumption (B7), we have that for $j\in\mathcal{A}_{>}$,
$G_{\gamma}(\cdot,\xi;j)^{\prime}(\bar{\theta};\bar{v})=\psi_{j}^{\prime}(f(\bar{\theta},\xi);f(\cdot,\xi)^{\prime}(\bar{\theta};\bar{v}))=\psi_{j}^{\prime}\left(\bar{z};-\frac{\bar{z}}{\left\lVert\bar{\theta}\right\rVert}\right)\leq-\frac{c_{\psi}l_{\psi}}{\left\lVert\bar{\theta}\right\rVert}\leq-\frac{c_{\psi}l_{\psi}}{\sqrt{q}\mu},$
@@
$\min_{v\in\mathcal{T}(\bar{\theta};\Theta);\left\lVert v\right\rVert\leq 1}G_{\gamma}^{\prime}(\bar{\theta};v)\leq G_{\gamma}^{\prime}(\bar{\theta};\bar{v})\leq\frac{1}{n}G_{\gamma}(\cdot,\bar{\xi};\bar{j})^{\prime}(\bar{\theta};\bar{v})\leq-\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu},$
@@
Finally, from Proposition 9.6 and convergence results in Sections 3.2.3 and 8.4, we summarize the convergence properties of the ESMM with exact penalization in the theorem below, where Assumptions (B1) to (B5) are adjusted to surrogate families of $F(\cdot,\xi)$ and $\psi_{j}(f(\cdot,\xi))$ for any $\xi\in\Xi$ and $j\in[J]$, and Assumption (B6) is adjusted to $V(\cdot;\lambda)$. The proof is similar to those of Theorems 3.6-8.12, and thus is omitted here.
Theorem 9.8 (Convergence of ESMM with exact penalization)
Under Assumptions (B1) to (B4) and the same settings of Proposition 9.6, for problem (9.2) with $\lambda\geq\bar{\lambda}$, let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon\geq 0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$. The following statements hold: